# Ginny Wong, Simon See

### v1

@@ 1 Introduction
The study of the minimal embeddable dimension is fundamentally related to several classic research directions, including the VC-dimension in statistical learning theory (Mohri et al., 2018) and the $k$-set problem in combinatorial geometry (Matousek, 2013). It is widely acknowledged that this question is related to the fundamental boundary underlying modern information retrieval (Lee et al., 2019; Weller et al., 2025b), data compression (Andoni and Indyk, 2008), and representation learning (Izacard et al., 2021; Wang et al., 2022).
@@
This paper questions the conclusions by Weller et al. (2025a) by studying the MED problem, as will be formalized in Section 2, from the perspective of approximability. That means we also work on the fundamental property of geometric spaces as in Weller et al. (2025a) but with an extended set of essential scoring functions: in addition to the inner product previously discussed (Weller et al., 2025a), cosine similarity and Euclidean distances are also included due to their wide application in embedding-based systems and representation learning. This paper discusses when those spaces are large enough to contain the subset membership of at most $k$ elements. This perspective, due to its theoretical nature, is not practical enough to help develop scalable ML algorithms. However, it still allows people to determine whether the effectiveness of embedding-based systems is limited by approximability (a fundamental property of geometric spaces) or by learnability (the way we build specific retrieval systems). If it turns out to be an approximability issue, any methodological effort might make little progress before the dimension of the vector space is large enough. Otherwise, there is still hope in developing more performant embedding-based retrieval systems.
Theoretical Contribution. We establish a clear quantitative relation $d=\Theta(k)$ between the MED $d$ and the maximum cardinality $k$ of the answer set for all three scoring functions. The major results are presented in Table 1, “Standard setting (MED)”. Our theoretical results are independent of the universe’s size $m$, suggesting that the fundamental property of geometric space does not limit the embedding-based retrieval system. They sharpen the theoretical bounds by Weller et al. (2025a) and contradict the empirical results in Weller et al. (2025a).
@@
In our simulation, we further assume the query vector of $q$ to be given by ${\bm{w}}_{q}:=\frac{1}{|S|}\sum_{x\in S}{\bm{x}}$ as the centroid of vectors in the answer set $S$ and denoted as the centroid setting. Our centroid setting is “more achievable” than the “free embedding optimization” setting (Weller et al., 2025a) because our setting only requires optimising $m$ embeddings for elements, but requires no optimization of the $\binom{m}{k}$ embeddings of the subsets. Simulations in both settings only show the existence of an embeddable configuration of $m$ elements in specific dimensions. By definition, they reveal upper bounds on MED.
Another interesting finding is that the simulation in the centroid setting reveals an empirical logarithmic upper bound of MED. This $O(\log m)$ dependency agrees with our theoretical upper bounds from a probabilistic construction. The minimal embeddable dimension in the centroid setting is denoted MED-C, and listed in Table 1. Compared to the polynomial dependency revealed by the free embedding optimization (Weller et al., 2025a), it might be counterintuitive because the centroid setting has fewer degrees of freedom ($\binom{m}{k}$ subset query embeddings are not optimized in the centroid setting), providing a perfect example of the bottleneck in learnability rather than approximability. More discussion can be found in Section 4.
@@ 2.1 $k$-shatter problem
$\textsc{MED}(m,k-1;\mathcal{F})$ $\leq\textsc{MED}(m,k;\mathcal{F})$
$\leq\textsc{MED}(m+1,k;\mathcal{F}).$
@@ 2.2 General bounds of MED by VC dimension
The definition of $k$-shattering also defines the VC dimension (Mohri et al., 2018), which is redefined below in terms of $k$-shattering.
@@
VC dimension is a measure of the capacity of classes of sets or binary functions (Vapnik, 2013). This concept plays a fundamental role in statistical learning theory, particularly in the study of approximability. By defining both the MED and VC dimensions using $k$-shattering, it is reasonable to expect some connections to hold between them, and indeed they do. For notational convenience, we define the “inversion" of VC dimension as
$\textsc{VCD}^{-1}(m;\mathcal{F})=\min\{i\mid s.t.\textsc{VCD}(i;\mathcal{F})\geq m\}$
,namely the minimum dimension required. Notably, the function $\textsc{VCD}(;\mathcal{F})$ is not guaranteed to increase strictly; therefore, there is no inverse function formally.
@@
Let $n=\textsc{VCD}^{-1}(m,\mathcal{F})$, then (1) $m$ points in n can be $m$-shattered by $\mathcal{F}$ but (2) $m$ points in $^{(}n-1)$ cannot be $m$-shattered by $\mathcal{F}$. Those two claims imply, in the sense of MED, that (1) $\textsc{MED}(m,m,\mathcal{F})\leq n$ and (2) $\textsc{MED}(m,m,\mathcal{F})>n-1$. Thus $\textsc{MED}(m,m,\mathcal{F})=n$
@@ 2.3 MED in the centroid setting
MED-C and MED. In the original definition of MED, only the existence of functionals is required to prove the $k$-shattering definition. In the centroid setting, $k$-centroid shattering fixes the subset query embeddings ${\bm{c}}_{S}$, which further determine functionals explicitly. Noticing that the number of “free” functionals involved in $k$-shattering is $\binom{m}{k}$, the $k$-centroid setting reduces the freedom into as few as $m$ vectors, which facilitates numerical simulation. For the same reason, however, MED lower bounds MED-C because of the flexibility to freely determine $f(\cdot)$ other than the centroid one $s(\cdot,{\bm{c}}_{S})$.. Formally, we have the following proposition.
@@
MED-C and neural set embeddings. One of the common deep learning approaches to compute a set embedding is to aggregate the embeddings of contained elements (Zaheer et al., 2017). We can also study the Minimum Embeddable Dimension in the Neural set embedding setting (MED-N), where the set embedding ${\bm{w}}_{S}={\rm NN}(\{{\bm{x}}:x\in S\})$ is computed by a complex neural architecture (such as MLP or multi-head self attention). The centroid setting simplifies the aggregation of complex neural networks by simple averaging. Therefore, one could expect that the MED-C is not smaller than the MED-N due to a less capable averaging operation. Combined with the definition of MED, this informal discussion suggested that we can roughly consider MED-N to be between MED and MED-C.
@@ 3.1 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\rm linear})$
To begin with, let’s consider the cyclic polytope (Ziegler, 2012).
@@
Given moment curve ${\bm{x}}(t)=(1,t,t^{2},...,t^{d})\in^{d},0\leq t\leq 1$, a cyclic polytope is the convex hull ${\rm Conv}({\bm{x}}(t_{1}),{\bm{x}}(t_{2}),\dots,{\bm{x}}(t_{m}))$, where $t_{i},i=1,...,m$ can be arbitrary real number.
One of the most well-known results of cyclic polytopes is that a cyclic polytope in d is an $\lfloor\frac{d}{2}\rfloor$-neighborly polytope, in which every $k$ vertices form a face when $k\leq\lfloor\frac{d}{2}\rfloor$ (Ziegler, 2012). By forming a $k$-face, it means that the $k$ vertices in this face can be linearly separated from the rest of the vertices of the polytope.
@@
Noticing the VC dimension of $\mathcal{F}_{\rm linear}$ in n is $n+1$ (Mohri et al., 2018), the lower bound is derived as a direct result of Proposition 2.8. The upper bound is by the cyclic polytope construct in Example 3.1. ∎
@@ 3.3 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\cos})$
$\textsc{MED}(m,k;\mathcal{F}_{\rm linear}))$ $\leq\textsc{MED}(m,k;\mathcal{F}_{\cos})$
$\leq\textsc{MED}(m,k;\mathcal{F}_{\rm linear}))+1.$
@@ 3.4 Discussion on $\Theta(k)$ bound of MEDs
The real performance bottleneck of the embedding-based retrieval. The bounds clearly show that the geometric space does not limit the effectiveness of the embedding-based retrieval system. To make this even more transparent, one possible strategy to build a perfect top-$k$ queries of a retrieval system with $m$ elements is to (1) put elements in the vertices of a cyclic polytope in 2k, (2) learn the query embeddings model which predict the query embedding that clearly separates the (at most $k$) answers from the rest. Our theory guarantees that such query embeddings exist, but the real problem is how to build a function to predict them. Knowing that neural networks can universally approximate functions (Hornik et al., 1989), the bottleneck of embedding-based retrieval is essentially a learnability problem: how to actually learn such a function with embedding dimension $2k$ from data.
@@ 4.1 Upper bounds
Although the $O(k^{2}\log m)$ bound may not be tight, it still suffices to make significant implications. An easy lower bound for MED-C is the lower bound of MED, which is $\Omega(k)$. We can see that the gap between the lower and upper bounds is a factor of $O(\log m)\times O(k)$. We discuss those two factors separately. For the factor of $O(\log m)$, it still demonstrates that a sublinear growth for MED-C with the size $m$ of the universe is feasible. For more complex neural set embedding settings, such as MED-N discussed in Section 2.3, a similar upper bound also applies. For the additional factor $O(k)$, we would like to convince the readers that $O(k\log m)$ is not very easy to achieve, so the $O(k^{2}\log m)$ is not very bad. To the best of the authors’ knowledge, the most similar (but not exact) problem setting to MED-C is 1-bit compress sensing (Aksoylar and Saligrama, 2014), with a lower bound of $\Omega(k\log\frac{m}{k})$, which cancels the additional $O(k)$ factor (compare $O(k^{2}\log m)$ for MED-C with $\Theta(k)$ for MED) but requires an additional sparse recovery algorithm rather than simply comparing scores independently. This related work suggests that the extra cost should be paid to achieve the bound $O(k\log m)$ of MED-C.
@@ 4.2 Numerical simulation
Experiment settings. We use optimize $m$ embeddings randomly initialized in by standard normal distribution in d. The objective function is a hinge loss computed over all pairwise positive and negative pairs across the top-$k$ queries. We use the Adam optimizer (Kingma, 2014) with a learning rate of 1. The optimization stops either after 1,000 steps or after every query finds its perfect answers. To facilitate full reproducibility, we list our Python code in the Appendix to disclose other details.
Comparison with baselines Previous work (Weller et al., 2025a) didn’t provide a quantitative relation in their theory, but instead offered an empirical relation via optimization of free embeddings, which corresponds to the standard setting in this paper. They search for the critical cardinality of the universe $m^{*}$ in a fixed dimension $d$ for the $k=2$ case, and established the following empirical results. We list their curve in Equation 1.
@@
It is a crucial question: how can the critical points identified by simulation in the centroid setting be compared with those found in the free embedding optimization setting (Weller et al., 2025a). The goal of numerical simulation is to estimate the minimum dimension such that $m$ elements can be $k$-shattered in some sense (or the maximum number of elements can be located in a given dimension). Therefore, all critical points found by any numerical simulations only suggest an upper bound of the minimal dimension (or lower bound of the maximum number of elements). The goal of this paper is to show that the MED (or the lower bound) grows slowly enough so that the fundamental property of the geometric space does not bottleneck the embedding-based retrieval. Therefore, results from simulations, whether computed from a centroid setting or a free embedding optimization setting, can serve as empirical evidence of the MED upper bound.
To compare with their empirical results, we also ran the experiments with $k=2$ on the centroid setting. We plot the results in Figure 1(a) and Figure 1(b). Figure 1(a) suggests that the number of critical points found in our centroid setting surpasses the curve fitted by Weller et al. (2025a) easily. From Fig 1(b), we can see from the results that our theory of MED-C agrees well with the empirical log-linear fitting, which is of a lower complexity family than the fitted result in Weller et al. (2025a).
@@ 4.3 Practical implication
Regarding the bottleneck of the embedding-based retrieval. The empirical results above suggested that the actual limitation of embedding-based retrieval, i.e., the number of objects that can be embedded, actually grows at least exponentially with the dimension. This contradicts the previous claims made by Weller et al. (2025a): The stagnation of the performance in the embedding-based system does NOT necessarily come from the fundamental geometric property of the vector space. The empirical findings resonated with the conclusion from $\Theta(k)$ bounds in a weaker form: we can observe the exponential growth of critical points and logarithmic growth of MED, but not a vertical line or horizontal line.
@@
$\overbrace{\min_{{\bm{w}}_{S}:S\in\mathcal{C}_{k}}}^{\text{additional variables}}\min_{{\bm{x}}:x\in X}L_{\rm FreeEmbedding}\leq\min_{{\bm{x}}:x\in X}L_{\rm Centroid},$ (2)
@@ 5 Related Theoretical Works
Guo et al. (2019) established the dependency of dimension $d$ and $K$ (the total number of queries) and $s$ (maximal number of queries associated with each object), as $d=O(s^{2}\log(K))$ in their Theorem 2.1. Applying to our study of MED, the total number of top $k$ queries of $m$ objects is $K:=\binom{m}{k}$, and the maximal number of queries on one object is $s:=\binom{m-1}{k-1}$. Plugging those two values back suggests an upper bound of MED by $O(\binom{m-1}{k-1}^{2}\log\binom{m}{k})$, which is worse than our $\Theta(k)$ bound.
You et al. (2025) studies the embedding-based retrieval in a specified "Hierarchical Retrieval (HR)" setting. It assumes a directed acyclic graph $\mathcal{G}$ between objects. The answer set $S(q_{i})$ of one query $q_{i}$ is defined by all "reachable objects" in $\mathcal{G}$ from the exact match object $E(q_{i})$. The upper bound is established as $d=O(\max(s\log m,\frac{1}{\epsilon^{2}}\log m)$, where $s$ is the largest cardinality of all answer sets. Applying their result in our top-$k$ query case, the graph $G$ is a $k$-clique. It still suggested a $\max(O(k\log m),O(\frac{1}{\epsilon^{2}})\log m)$ upper bound of MED, which is worse than our $\Theta(k)$ bound.

### v2

@@ 1 Introduction
The study of the minimal embeddable dimension is fundamentally related to several classic research directions, including the VC-dimension in statistical learning theory [9] and the $k$-set problem in combinatorial geometry [8]. It is widely acknowledged that this question is related to the fundamental boundary underlying modern information retrieval [7, 14], data compression [2], and representation learning [5, 12].
@@
This paper questions the conclusions by [13] by studying the MED problem, as will be formalized in Section 2, from the perspective of approximability. That means we also work on the fundamental property of geometric spaces as in [13] but with an extended set of essential scoring functions: in addition to the inner product previously discussed [13], cosine similarity and Euclidean distances are also included due to their wide application in embedding-based systems and representation learning. This paper discusses when those spaces are large enough to contain the subset membership of at most $k$ elements. This perspective, due to its theoretical nature, is not practical enough to help develop scalable ML algorithms. However, it still allows people to determine whether the effectiveness of embedding-based systems is limited by approximability (a fundamental property of geometric spaces) or by learnability (the way we build specific retrieval systems). If it turns out to be an approximability issue, any methodological effort might make little progress before the dimension of the vector space is large enough. Otherwise, there is still hope in developing more performant embedding-based retrieval systems.
Theoretical Contribution. We establish a clear quantitative relation $d=\Theta(k)$ between the MED $d$ and the maximum cardinality $k$ of the answer set for all three scoring functions. The major results are presented in Table 1, “Standard setting (MED)”. Our theoretical results are independent of the universe’s size $m$, suggesting that the fundamental property of geometric space does not limit the embedding-based retrieval system. They sharpen the theoretical bounds by [13] and contradict the empirical results in [13].
@@
In our simulation, we further assume the query vector of $q$ to be given by ${\bm{w}}_{q}:=\frac{1}{|S|}\sum_{x\in S}{\bm{x}}$ as the centroid of vectors in the answer set $S$ and denoted as the centroid setting. Our centroid setting is “more achievable” than the “free embedding optimization” setting [13] because our setting only requires optimising $m$ embeddings for elements, but requires no optimization of the $\binom{m}{k}$ embeddings of the subsets. Simulations in both settings only show the existence of an embeddable configuration of $m$ elements in specific dimensions. By definition, they reveal upper bounds on MED.
Another interesting finding is that the simulation in the centroid setting reveals an empirical logarithmic upper bound of MED. This $O(\log m)$ dependency agrees with our theoretical upper bounds from a probabilistic construction. The minimal embeddable dimension in the centroid setting is denoted MED-C, and listed in Table 1. Compared to the polynomial dependency revealed by the free embedding optimization [13], it might be counterintuitive because the centroid setting has fewer degrees of freedom ($\binom{m}{k}$ subset query embeddings are not optimized in the centroid setting), providing a perfect example of the bottleneck in learnability rather than approximability. More discussion can be found in Section 4.
@@ 2.1 $k$-shatter problem
$\textsc{MED}(m,k-1;\mathcal{F})$ $\leq\textsc{MED}(m,k;\mathcal{F})\leq\textsc{MED}(m+1,k;\mathcal{F}).$
@@ 2.2 General bounds of MED by VC dimension
The definition of $k$-shattering also defines the VC dimension [9], which is redefined below in terms of $k$-shattering.
@@
VC dimension is a measure of the capacity of classes of sets or binary functions [11]. This concept plays a fundamental role in statistical learning theory, particularly in the study of approximability. By defining both the MED and VC dimensions using $k$-shattering, it is reasonable to expect some connections to hold between them, and indeed they do. For notational convenience, we define the “inversion" of VC dimension as
$\textsc{VCD}^{-1}(m;\mathcal{F})=\min\{i\mid s.t.\textsc{VCD}(i;\mathcal{F})\geq m\},$
namely the minimum dimension required. Notably, the function $\textsc{VCD}(\cdot;\mathcal{F})$ is not guaranteed to increase strictly; therefore, there is no inverse function formally.
@@
Let $n=\textsc{VCD}^{-1}(m,\mathcal{F})$, then (1) $m$ points in n can be $m$-shattered by $\mathcal{F}$ but (2) $m$ points in (n-1) cannot be $m$-shattered by $\mathcal{F}$. Those two claims imply, in the sense of MED, that (1) $\textsc{MED}(m,m,\mathcal{F})\leq n$ and (2) $\textsc{MED}(m,m,\mathcal{F})>n-1$. Thus $\textsc{MED}(m,m,\mathcal{F})=n$
@@ 2.3 MED in the centroid setting
MED-C and MED. In the original definition of MED, only the existence of functionals is required to prove the $k$-shattering definition. In the centroid setting, $k$-centroid shattering fixes the subset query embeddings ${\bm{c}}_{S}$, which further determine functionals explicitly. Noticing that the number of “free” functionals involved in $k$-shattering is $\binom{m}{k}$, the $k$-centroid setting reduces the freedom into as few as $m$ vectors, which facilitates numerical simulation. For the same reason, however, MED lower bounds MED-C because of the flexibility to freely determine $f(\cdot)$ other than the centroid one $s(\cdot,{\bm{c}}_{S})$. Formally, we have the following proposition.
@@
MED-C and neural set embeddings. One of the common deep learning approaches to compute a set embedding is to aggregate the embeddings of contained elements [17]. We can also study the Minimum Embeddable Dimension in the Neural set embedding setting (MED-N), where the set embedding ${\bm{w}}_{S}={\rm NN}(\{{\bm{x}}:x\in S\})$ is computed by a complex neural architecture (such as MLP or multi-head self attention). The centroid setting simplifies the aggregation of complex neural networks by simple averaging. Therefore, one could expect that the MED-C is not smaller than the MED-N due to a less capable averaging operation. Combined with the definition of MED, this informal discussion suggested that we can roughly consider MED-N to be between MED and MED-C.
@@ 3.1 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\rm linear})$
To begin with, let’s consider the cyclic polytope [18].
@@
Given moment curve ${\bm{x}}(t)=(1,t,t^{2},...,t^{d})\in^{d},0\leq t\leq 1$, a cyclic polytope is the convex hull ${\rm Conv}({\bm{x}}(t_{1}),{\bm{x}}(t_{2}),\dots,{\bm{x}}(t_{m}))$, where $t_{1}<t_{2}<\cdots<t_{m}$ can be arbitrary real numbers.
One of the most well-known results of cyclic polytopes is that a cyclic polytope in d is an $\lfloor\frac{d}{2}\rfloor$-neighborly polytope, in which every $k$ vertices form a face when $k\leq\lfloor\frac{d}{2}\rfloor$ [18]. By forming a $k$-face, it means that the $k$ vertices in this face can be linearly separated from the rest of the vertices of the polytope.
@@
Noticing the VC dimension of $\mathcal{F}_{\rm linear}$ in n is $n+1$ [9], the lower bound is derived as a direct result of Proposition 2.8. The upper bound is by the cyclic polytope construct in Example 3.1. ∎
@@ 3.3 Tight bounds of $\textsc{MED}(m,k;\mathcal{F}_{\cos})$
$\textsc{MED}(m,k;\mathcal{F}_{\rm linear}))$ $\leq\textsc{MED}(m,k;\mathcal{F}_{\cos})\leq\textsc{MED}(m,k;\mathcal{F}_{\rm linear}))+1.$
@@ 3.4 Discussion on $\Theta(k)$ bound of MEDs
The real performance bottleneck of the embedding-based retrieval. The bounds clearly show that the geometric space does not limit the effectiveness of the embedding-based retrieval system. To make this even more transparent, one possible strategy to build a perfect top-$k$ queries of a retrieval system with $m$ elements is to (1) put elements in the vertices of a cyclic polytope in 2k, (2) learn the query embeddings model which predict the query embedding that clearly separates the (at most $k$) answers from the rest. Our theory guarantees that such query embeddings exist, but the real problem is how to build a function to predict them. Knowing that neural networks can universally approximate functions [4], the bottleneck of embedding-based retrieval is essentially a learnability problem: how to actually learn such a function with embedding dimension $2k$ from data.
@@ 4.1 Upper bounds
Although the $O(k^{2}\log m)$ bound may not be tight, it still suffices to make significant implications. An easy lower bound for MED-C is the lower bound of MED, which is $\Omega(k)$. We can see that the gap between the lower and upper bounds is a factor of $O(\log m)\times O(k)$. We discuss those two factors separately. For the factor of $O(\log m)$, it still demonstrates that a sublinear growth for MED-C with the size $m$ of the universe is feasible. For more complex neural set embedding settings, such as MED-N discussed in Section 2.3, a similar upper bound also applies. For the additional factor $O(k)$, we would like to convince the readers that $O(k\log m)$ is not very easy to achieve, so the $O(k^{2}\log m)$ is not very bad. To the best of the authors’ knowledge, the most similar (but not exact) problem setting to MED-C is 1-bit compress sensing [1], with a lower bound of $\Omega(k\log\frac{m}{k})$, which cancels the additional $O(k)$ factor (compare $O(k^{2}\log m)$ for MED-C with $\Theta(k)$ for MED) but requires an additional sparse recovery algorithm rather than simply comparing scores independently. This related work suggests that the extra cost should be paid to achieve the bound $O(k\log m)$ of MED-C.
@@ 4.2 Numerical simulation
Experiment settings. We use optimize $m$ embeddings randomly initialized in by standard normal distribution in d. The objective function is a hinge loss computed over all pairwise positive and negative pairs across the top-$k$ queries. We use the Adam optimizer [6] with a learning rate of 1. The optimization stops either after 1,000 steps or after every query finds its perfect answers. To facilitate full reproducibility, we list our Python code in the Appendix to disclose other details.
Comparison with baselines Previous work [13] didn’t provide a quantitative relation in their theory, but instead offered an empirical relation via optimization of free embeddings, which corresponds to the standard setting in this paper. They search for the critical cardinality of the universe $m^{*}$ in a fixed dimension $d$ for the $k=2$ case, and established the following empirical results. We list their curve in Equation 1.
@@
It is a crucial question: how can the critical points identified by simulation in the centroid setting be compared with those found in the free embedding optimization setting [13]. The goal of numerical simulation is to estimate the minimum dimension such that $m$ elements can be $k$-shattered in some sense (or the maximum number of elements can be located in a given dimension). Therefore, all critical points found by any numerical simulations only suggest an upper bound of the minimal dimension (or lower bound of the maximum number of elements). The goal of this paper is to show that the MED (or the lower bound) grows slowly enough so that the fundamental property of the geometric space does not bottleneck the embedding-based retrieval. Therefore, results from simulations, whether computed from a centroid setting or a free embedding optimization setting, can serve as empirical evidence of the MED upper bound.
To compare with their empirical results, we also ran the experiments with $k=2$ on the centroid setting. We plot the results in Figure 1(a) and Figure 1(b). Figure 1(a) suggests that the number of critical points found in our centroid setting surpasses the curve fitted by [13] easily. From Fig 1(b), we can see from the results that our theory of MED-C agrees well with the empirical log-linear fitting, which is of a lower complexity family than the fitted result in [13].
@@ 4.3 Practical implication
Regarding the bottleneck of the embedding-based retrieval. The empirical results above suggested that the actual limitation of embedding-based retrieval, i.e., the number of objects that can be embedded, actually grows at least exponentially with the dimension. This contradicts the previous claims made by [13]: The stagnation of the performance in the embedding-based system does NOT necessarily come from the fundamental geometric property of the vector space. The empirical findings resonated with the conclusion from $\Theta(k)$ bounds in a weaker form: we can observe the exponential growth of critical points and logarithmic growth of MED, but not a vertical line or horizontal line.
@@
$\overbrace{\min_{{\bm{w}}_{S}:S\in\mathcal{C}_{k}}}^{\text{additional variables}}\min_{{\bm{x}}:x\in X}L_{\rm FreeEmbedding}\leq\min_{{\bm{x}}:x\in X}L_{\rm Centroid},$
@@ 5 Related Theoretical Works
[3] studies a multi-class classification problem over embeddings. Their work established the dependency of dimension $d$ and $K$ (the total number of classes) and $s$ (maximal number of classes associated with each object), as $d=O(s^{2}\log(K))$ in their Theorem 2.1. Applying to our study of MED, the total number of top $k$ queries of $m$ objects defines the total number of classes as $K:=\binom{m}{k}$, and the maximal number of classes on one object is $s:=\binom{m-1}{k-1}$. Plugging those two values back suggests an upper bound of MED by $O(\binom{m-1}{k-1}^{2}\log\binom{m}{k})$, which is significantly worse than our $\Theta(k)$ bound.
[16] studies the embedding-based retrieval in a specified "Hierarchical Retrieval (HR)" setting. It assumes a directed acyclic graph $\mathcal{G}$ between objects. The answer set $S(q_{i})$ of one query $q_{i}$ is defined by all "reachable objects" in $\mathcal{G}$ from the exact match object $E(q_{i})$, i.e., $E(q_{i})$ and all its neighbors in $\mathcal{G}$. The upper bound is established as $d=O(\max(s\log m,\frac{1}{\epsilon^{2}}\log m))$, where $s$ is the largest cardinality of all answer sets. However, we argue that the HR setting and our study of top-$k$ queries are fundamentally different, which is clearly revealed by considering the number of effective queries or the total number of different answer sets. In the HR setting, noticing that each answer set is defined by the reachability in a DAG, we can conclude that there are no more than $m$ different answer sets. In a sharp contrast, the total number of answer sets of all top-$k$ queries is $\binom{m}{k}\approx k^{m}$. Translating their results in such a strong assumption still suggests a $\max(O(k\log m),O(\frac{1}{\epsilon^{2}})\log m)$ upper bound of MED, which is worse than our $\Theta(k)$ bound due to its dependency on $\log m$.